OpenStudy (anonymous):

What is the 5th term of an arithmetic sequence if t2 = -5 and t6 = 7?

OpenStudy (anonymous):

\[t_6-t_2=(6-2)d\] where \(d\) is the common difference between consecutive terms. \[7-(-5)=4d~~\implies~~d=\frac{12}{4}=3\] You have \(t_6-t_5=d\), or \(7-3=t_5\).

OpenStudy (anonymous):

So the answer is 4?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you! <3